# From High-SNR Radar Signal to ECG: A Transfer Learning Model with Cardio-Focusing Algorithm for Scenarios with Limited Data

### v1

@@ V-B2 Evaluations of Fine-tuning Results
The fine-tuning is based on the model pre-trained by $100\%$ dataset with or without sparse penalty, and the experiment is repeated for different percentages of labeled data (i.e., radar signal with ECG ground truth). In addition, the same deep-learning model will be trained in a supervised manner by using the same amount of labeled data as the reference for transfer-learning ECG recovery. At last, an overall improvement will be provided by calculating the percentage of improvement across all four metrics to provide straightforward evaluations for different methods, as shown in Table IV.
The fine-tuning with $100\%$ labeled data provides very similar performance in morphological accuracy with the MSE and PCC around $0.0080$ and $85.47\%$. It is worth noticing that the peak error and MDR are slightly improved, because the pre-text task SSR for pre-training is equivalent to identifying the peak position of the radar signal, and the learned representations can be seemly transferred to improve the accuracy of the recovered ECG R peaks, contributing to the overall improvement for transfer learning ($3.66\%$ and $1.75\%$).
Reducing $20\%$ of labeled data causes a $10\%$ overall degradation as shown in Table IV, and the decline of peak error and MDR is more than MSE and PCC. The reason is that ECG morphological patterns for different cardiac cycles are similar and can be well-learned from $80\%$ labeled data with good MSE and PCC ($0.0084$ and $84.60\%$), while the location of each ECG piece is random and can be distorted by noises, requiring more training data for convergence.
The supervised training with $60\%$ labeled data cannot ensure a good morphological and peak accuracy and the overall degradation is $23.37\%$, with the PCC drop below $80\%$ as shown in Table IV. In contrast, the pre-trained model still provided good results with mild degradations of $8.71\%$ and $12.68\%$. It is noticed that the effectiveness of sparse penalty in the SSL stage also affects the fine-tuning stage, because both peak error and MDR for transfer learning with spare penalty are better than those without sparse penalty, causing a large gap in the overall improvement compared with the previous training with $100\%$ and $80\%$ dataset.
Lastly, the deep learning model can barely learn from $40\%$ labeled dataset and yield a bad morphological and peak accuracy for supervised learning. In addition, examples of the recovered ECG for transfer learning with or without sparse penalty are shown in Figure 11, and it is clear that the deep learning model struggles to learn both morphological and peak features from limited data if the pre-text task is not well-trained without sparse penalty, while Figure LABEL:sub@fig:fs_good shows the good recovery because the pre-trained model transfer the learned representations from radar inputs to ECG recovery.
@@
TABLE IV: Performance of ECG Recovery using different percentages of labeled data
Methods MSE (×$\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ Overall $\uparrow$     MSE (×$\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ Overall $\uparrow$
$\mathbf{100\%}$ Labeled     $\mathbf{80\%}$ Labeled
Supervised 0.800.80{0.80}0.80 ${85.47\%}$ 7.617.61{7.61}7.61 ${6.85\%}$ $0.00\%$     $0.84$ $84.60\%$ $8.90$ $8.04\%$ $-10.09\%$
TF w/o sp $0.81$ $85.35\%$ $8.46$ $5.51\%$ $1.75\%$     $0.82$ $86.36\%$ $8.35$ $7.02\%$ $-3.42\%$
TF with sp $0.80$ $85.51\%$ $8.40$ $5.14\%$ $3.66\%$     $0.81$ $84.29\%$ $8.31$ $7.14\%$ $-4.02\%$
$\mathbf{60\%}$ Labeled     $\mathbf{40\%}$ Labeled
Supervised $0.93$ ${79.91\%}$ 10.6510.65{10.65}10.65 ${8.93\%}$ $-23.27\%$     $0.98$ $75.89\%$ $11.15$ $12.15\%$ $-39.4\%$
TF w/o sp $0.85$ $83.74\%$ $8.84$ $8.65\%$ $-12.68\%$     $0.97$ $76.56\%$ $10.87$ $10.99\%$ $-34.05\%$
TF with sp $0.86$ $84.92\%$ $8.58$ $7.72\%$ $-8.71\%$     $0.93$ $78.72\%$ $8.70$ $9.02\%$ $-17.54\%$
TF for transfer learning and sp for sparse penalty
@@
Figure 11: Results of transfer learning using limited labeled data: (a) Poor ECG recovery without proper morphological feature and peak location; (b) Good ECG recovery owing to the pre-trained model.
V-B3 Summary of Transfer-learning-based ECG Recovery
@@
Previous evaluations in terms of SSL and fine-tuning stages have illustrated the ability of the proposed RFcardi to learn from unlabeled data and transfer the knowledge to the ECG recovery task using limited radar-ECG pairs. The overall performance in Table IV are plotted in Figure 12 for a straightforward comparison:
@@
  • Transfer learning could enhance the performance of ECG recovery for the cases with ample labeled training data, and the quality of the pre-trained model has a minor effect on the final result because the deep learning model could learn from numerous radar-ECG pairs.
  • For the cases with limited labeled data ($40\%$, $60\%$), the proposed RFcardi shows outstanding performance owing to the representations learned from unlabeled data. Furthermore, the quality of the pre-trained model does matter to alleviate the burden of deep learning model to learn both morphological ECG patterns and peak locations, as indicated by the increasing gap between orange and green lines in Figure 12.
Figure 12: Overall performance of the radar-based ECG recovery.
@@ VI Conclusions
This paper investigates the efficient collection of high-SNR radar signals with ample cardiac features for transfer-learning-based ECG recovery. Previous methods adopted signal accumulation or clustering to suppress the noises, while the rough localization based on FMCW radar cannot accurately reveal the chest region, requiring a time-consuming traverse among a 3D space for compensation. In this paper, a novel CFT algorithm is proposed to dynamically articulate the points with the best SNR and could track the cardiac location over time if the subjects change posture. In addition, a transfer learning framework RFcardi is designed with SSR as a pre-text task for pre-training to reduce the dependency on cumbersome ECG ground truth collection. The experiments performed in different scenarios prove the feasibility of the CFT-RFcardi framework in radar signal extraction and ECG recovery with limited labeled data, enabling a convenient deployment in new scenarios with limited data for future contactless wellness monitoring.

### v2

@@ V-B2 Evaluations and Ablation Studies of Fine-tuning Results
To verify the impact of SSL quality on the fine-tuning result, the same amount of labeled data is used for fine-tuning with different sparse penalty $\lambda_{s}$, as shown in Table IV. The overall improvements are also provided in the last column by calculating the percentage of improvement across all four metrics to provide straightforward evaluations for different methods, as shown in Table IV and Figure LABEL:sub@fig:rf_overall.
The results coincide with the expectation that more labeled data generates better ECG recovery performance, and the discrepancy between using different $\lambda_{s}$ is not obvious if the training data is enough (e.g., for $100\%$ and $80\%$), as indicated in Figure LABEL:sub@fig:rf_overall. However, if the labeled data is not enough, the best pre-training model ($\lambda_{s}=0.01$) gives the best ECG recovery for all cases with different labeled data scales, because SSL with $\lambda_{s}=0.02$ or 0 does not properly train the backbone with the ability to capture the peak-related information. In addition, the gap between using different $\lambda_{s}$ is larger than the case with enough data as shown in Figure LABEL:sub@fig:rf_overall, revealing the advantage of using SSL for pre-training. According to our experiment, $\lambda_{s}=0.01$ guarantees a good performance for different scales of input datasets and will be adopted for the comparison with other SSL methods later.
Figure 12: Overall performance of the radar-based ECG recovery: (a) RFcardi with different $\lambda_{s}$; (b) Comparison with different SSL methods.
@@
TABLE V: Performance of ECG recovery using different percentages of labeled data
Methods MSE ($\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ Overall $\uparrow$     MSE ($\times 10^{-2}$) $\downarrow$ PCC $\uparrow$ Peak Error (ms) $\downarrow$ MDR $\downarrow$ Overall $\uparrow$
$\mathbf{100\%}$ Unlabeled, $\mathbf{100\%}$ Labeled     $\mathbf{100\%}$ Unlabeled, $\mathbf{80\%}$ Labeled
Supervised 0.80{0.80} $85.47\%$} 7.61{7.61} $6.85\%$} $0.00\%$     $0.84$ $84.60\%$ $8.90$ $8.04\%$ $10.09\%$
MaeFE [33] $0.81$ $83.97\%$ $7.98$ $5.89\%$ $1.54\%$     $0.84$ $83.61\%$ $9.22$ $6.86\%$ $7.12\%$
CL [34] $0.81$ $6.15\% 8.8$58.85 $4.47\%$ $3.22\%$}     $0.82$ $84.30\%$ $9.96$ $6.00\%$ $5.59\%$
RFcardi w/o sp $0.81$ $85.35\%$ $8.46$ $5.51\%$ $1.75\%$     $0.82$ $86.36\%$ $8.35$ $7.02\%$ $\mathbf{-3.42\%}$
RFcardi with sp $0.80$ $85.51\%$ $8.40$ $5.14\%$ $\mathbf{3.66\%}$     $0.81$ $84.29\%$ $8.31$ $7.14\%$ $4.02\%$}
$\mathbf{100\%}$ Unlabeled, $\mathbf{60\%}$ Labeled     $\mathbf{100\%}$ Unlabeled, $\mathbf{40\%}$ Labeled
Supervised $0.93$ $79.91\%$} 10.65{10.65} $8.93\%$} $23.27\%$     $0.98$ $75.89\%$ $11.15$ $12.15\%$ $39.40\%$
MaeFE [33] $0.87$ $82.01\%$ $9.46$ $8.68\%$ $15.96\%$     $0.96$ $78.51\%$ $10.36$ $8.51\%$ $22.13\%$}
CL [34] $0.92$ $78.83\%$ $10.12$ $12.42\%$ $21.65\%$     $1.05$ $75.39\%$ $11.03$ $11.25\%$ $38.05\%$
RFcardi w/o sp $0.85$ $3.74\% 8.8$48.84 $8.65\%$ $12.68\%$}     $0.97$ $76.56\%$ $10.87$ $10.99\%$ $34.05\%$
RFcardi with sp $0.86$ $84.92\%$ $8.58$ $7.72\%$ $\mathbf{-8.71\%}$     $0.93$ $78.72\%$ $8.70$ $9.02\%$ $\mathbf{-17.54\%}$
sp for sparse penalty ($\lambda_{s}=0.01$)
V-B3 Comparison with Other SSL Methods
The comparison is based on the model pre-trained on a $100\%$ dataset using different SSL methods, and the experiment is repeated for different percentages of labeled data (i.e., radar signal with ECG ground truth). In addition, the same deep-learning model will be trained in a supervised manner by using the same amount of labeled data as the benchmark for transfer-learning ECG recovery. The overall performances are also listed in the last column and plotted in Figure LABEL:sub@fig:overall.
The fine-tuning with $100\%$ labeled data provides very similar performance in morphological accuracy, with the MSE and PCC around $0.0080$ and $85.47\%$. It is worth noticing that the peak error and MDR are slightly improved, because the pre-text task SSR for pre-training is equivalent to identifying the peak position of the radar signal, and the learned representations can be seamlessly transferred to improve the accuracy of the recovered ECG R peaks, contributing to the overall improvement for RFcardi ($3.66\%$ and $1.75\%$). Similarly, MaeFE and CL also achieve slight improvement, although they are not pre-trained by the cardiac-feature-related pre-text tasks, enough data still guarantees an effective pre-training and fine-tuning.
Reducing $20\%$ of labeled data causes a $10\%$ overall degradation for supervised learning as shown in Table V, and the decline of peak error and MDR is more than MSE and PCC. The reason is that ECG morphological patterns for different cardiac cycles are similar and can be well-learned from $80\%$ labeled data with good MSE and PCC ($0.0084$ and $84.60\%$), while the location of each ECG piece is random and can be distorted by noise, requiring more training data for convergence. In addition, $80\%$ dataset is still enough for MaeFE and CL to achieve a similar performance with RFcardi, with the performance of $7.12\%$ and $5.19\%$ compared with $4.02\%$ for RFcardi with sparse penalty.
The supervised training with $60\%$ labeled data cannot ensure a good morphological and peak accuracy, and the overall degradation is $23.37\%$, with the PCC drop below $80\%$ as shown in Table V. In contrast, the pre-trained model still provided good results with mild degradations of $8.71\%$ and $12.68\%$. It is noticed that the effectiveness of the sparse penalty in the SSL stage also affects the fine-tuning stage, because both peak error and MDR for transfer learning with sparse penalty are better than those without sparse penalty, causing a large gap in the overall improvement compared with the previous training with $100\%$ and $80\%$ datasets. In this stage, CL shows a larger gap compared with other SSL methods as shown in Figure LABEL:sub@fig:overall, with the overall performance of $21.65\%$, close to the result achieved by supervised learning ($23.27\%$). The reason is that CL relies on the distinction between positive and negative samples, while the construction designed in [34] is not suitable for cardiac signals, because the cardiac activities may share the same pattern even after applying more than $20$ seconds of delay.
Lastly, the deep learning model can barely learn from $40\%$ labeled dataset and yield a bad morphological and peak accuracy for supervised learning. In addition, CL, MaeFE, and RFcardi, without a sparse penalty, all struggle to learn both morphological and peak features from limited data if the pre-text task is not well-trained. CL and MaeFE do not leverage specially designed pre-text to highlight the peak-related features and cannot achieve an effective SSL with limited data, as evaluated in Section V-B2. Similarly, RFcardi without a sparse penalty cannot achieve convergence during SSL, with the examples of the recovered ECG signals shown in Figure 13. It is obvious that the well-pretrained model generates the ECG signal with good peak and morphological accuracy after fine-tuning as shown in Figure LABEL:sub@fig:fs_good, while the large peak deviation and bad ECG morphology are caused by the pre-trained models that cannot capture essential peak-related features, as shown in Figure LABEL:sub@fig:fs_bad.
@@ V-B3 Comparison with Other SSL Methods
Figure 13: Results of transfer learning using limited labeled data: (a) Poor ECG recovery without proper morphological feature and peak location; (b) Good ECG recovery owing to the well pre-trained model.
V-B4 Summary of Transfer-learning-based ECG Recovery
@@
Previous evaluations in terms of SSL and fine-tuning stages have illustrated the ability of the proposed RFcardi to learn from unlabeled data and transfer the knowledge to the ECG recovery task using limited radar-ECG pairs. Here are some key conclusions drawn from Figure 12:
@@
  • Transfer learning could enhance the performance of ECG recovery for cases with ample labeled training data, and the quality of the pre-trained model has a minor effect on the final result because the deep learning model could learn from numerous radar-ECG pairs.
  • The proposed RFcardi shows outstanding performance in learning latent representations from unlabeled data through a specially designed pre-text task, SSR. In comparison, MaeFE and CL are not designed for this task and cannot effectively capture the peak-related features when the input data is limited.
  • For the cases with limited labeled data ($40\%$, $60\%$), the quality of the pre-trained model does matter to alleviate the burden of the deep learning model to learn both morphological ECG patterns and peak locations, as indicated by the increasing gap in both Figure LABEL:sub@fig:rf_overall and LABEL:sub@fig:overall.
V-C Limitations and Future Work
V-C1 Real-time Processing
Figure 9 indicates that the time consumed for each $1$-minute trial is around $30$ sec, but the proposed CFT algorithm still needs improvement to achieve real-time processing, because the searching time can reach $6.5$ sec ($k_{max}=100$) for a $4$-sec segment if the collected raw radar signal is heavily distorted or the CF point is far from the current $E_{k}$. In future work, real-time processing ability should be considered as an essential ability, and the high-SNR signal collection algorithm should be tested on the edge devices to enable real-life applications.
V-C2 Energy Consumption
In this study, the CFT algorithm requires multiple function evaluations, and the RFcardi model involves matrix multiplications and convolutional operations that are computationally intensive during inference. Considering that the mobile platforms typically face strict power budgets, the searching strategy of CFT can be optimized to reduce point evaluations by using adaptive grid adjustment and search directions [35]. For deep learning, model pruning and knowledge distillation can decrease the computational load, reducing both runtime and energy demand during inference [36].
V-C3 Multi-target Vital Sign Monitoring
In real-world monitoring, a multi-target scenario will introduce extra interference and also require the separation of the signals reflected from different subjects [2]. The future work should focus on high-resolution and adaptive beamforming to better isolate the signal reflected from individual targets, and the CFT algorithm should be advanced to support multi-target vital signal extraction.
@@ VI Conclusions
This paper investigates the efficient collection of high-SNR radar signals with ample cardiac features for transfer-learning-based ECG recovery. Previous methods adopted signal accumulation or clustering to suppress the noise, while the rough localization based on FMCW radar cannot accurately reveal the chest region, requiring a time-consuming traverse through a 3D space for compensation. In this paper, a novel CFT algorithm is proposed to dynamically find the points with the best SNR and could track the cardiac location over time if the subjects change posture. In addition, a transfer learning framework RFcardi is designed with SSR as a pre-text task for pre-training to reduce the dependency on cumbersome ECG ground truth collection. The experiments performed in different scenarios prove the feasibility of the CFT-RFcardi framework in radar signal extraction and ECG recovery with limited labeled data, enabling a convenient deployment in new scenarios with limited data for remote healthcare monitoring.